OpenStudy (anonymous):

PLEASE HELP! Find the surface area of a cylinder in terms of pi with a radius of 3 and height of 6. Will fan and give medal to those who help!

OpenStudy (mathstudent55):

The surface area of a cylinder consists of three parts. The top and bottom bases which a circles. The side part, the lateral area, is a rectangle whose sides are the height of the cylinder and the circumference of the base.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Alright but I've used the formula and everything and still wrong.

OpenStudy (mathstudent55):

Area of a circle: \(\large A_{circle} = \pi r^2\) Area of a rectangle: \(\large A_{rectangle} = LW\) In this case, the rectangle has two side which are: one side is the circumference of the base: \(L = 2 \pi r\) the other side is the height of the cylinder: \(W = h\)

OpenStudy (mathstudent55):

That means the complete formula for the surface area of a cylinder is: \(A_{cylinder}= 2 \pi r^2 + 2\pi r h\)

OpenStudy (anonymous):

2*3.14*3^2+2*3.14*3*6 Is that right?

OpenStudy (mathstudent55):

Now use r = 3 and h = 6, an tell me what you get for the area. Remember the instructions tell you to leave the area in terms of pi. That means to leave pi in the result and to not use 3.14 as the value of pi.

OpenStudy (anonymous):

I get 169.56.

OpenStudy (mathstudent55):

\(\large A_{cylinder}= 2 \pi r^2 + 2\pi r h\) \(\large A_{cylinder}= 2 \times \pi \times 3^2 + 2\pi \times 3 \times 6\) You see how I replaced r and h with the given values, but left pi in the formula?

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

\(\large A_{cylinder}= 2 \times 3^2 \times + 2 \times 3 \times 6 \times \pi\)

OpenStudy (mathstudent55):

Multiply all numbers together and leave pi.

OpenStudy (mathstudent55):

\(\large A_{cylinder}= 2 \times 3^2 \times \pi+ 2 \times 3 \times 6 \times \pi\)

OpenStudy (anonymous):

131.0973355

OpenStudy (mathstudent55):

You are still multiplying by 3.14. All the numbers you have are whole numbers. You can't get any decimals.

OpenStudy (anonymous):

okay. sorry if i sound stupid.

OpenStudy (anonymous):

54?

OpenStudy (mathstudent55):

\(\large A_{cylinder}= 2 \times 3^2 \times \pi+ 2 \times 3 \times 6 \times \pi\) \(\large A_{cylinder}= 2 \times 9 \times \pi+ 36 \times \pi\) \(\large A_{cylinder}= 18 \times \pi+ 36 \times \pi\) \(\large A_{cylinder}= 18 \pi+ 36 \pi\) Since both terms have pi, they are like terms and can be combined, so you get: \(\large A_{cylinder}= 54 \pi\)

OpenStudy (mathstudent55):

Yes, the number is 54, but 54 is multiplying pi.

OpenStudy (mathstudent55):

Your answer above 169.56 was correct for an approximation, but they want an exact answer in terms of pi, so the answer is 54pi

OpenStudy (anonymous):

Okay thank you so much! ^^

OpenStudy (mathstudent55):

You're welcome.